# A simplified directional KeRF algorithm

### v1

@@ 2.2. The Centered KeRF algorithm
A simple observation from the theorem above is, that the infinite centered KeRF coincides with the infinite simplified directional KeRF. Hence, under specific assumptions, we can compute the rate of convergence of the simplified KeRF.
Corollary 1.
([IA24] Theorem 1.) Under the following assumptions:
$Y=m(X)+\epsilon$
$X\text{ is uniformly distributed on }[0,1]^{d}$
$\epsilon\sim\mathcal{N}(0,\sigma^{2}),\quad\sigma<\infty$
[…]
3. Plots and experiments.
In this final section, we conduct numerical simulations and experiments to compare the performance of the finite-centered KeRF algorithm and the finite-simplified directional KeRF algorithm. The evaluation is carried out in terms of the $L_{2}$-error and the standard deviation of the error for various target functions $Y$. Specifically, we generated a two-dimensional feature space of size $n=1500$ comprising uniformly distributed points.
  1. (1)
  2. (2)
  3. (3)
    $Y=2X_{1}+e^{-X_{2}^{2}}.$
(a) $L-$error for the function $Y=2X_{1}+e^{-X^{2}_{2}}$.
Figure 5. Comparison of $L-$error and standard deviation.
Figure 6. Comparison of $L-$error and standard deviation.
Figure 7. Comparison of $L-$error and standard deviation.
As one might expect, all the figures 5(a) 6(a), 7(a) exhibit similar behavior. For small values of trees, the two algorithms demonstrate slightly different performances, however as the number of trees increases consistent with the theorem 1, both algorithms have the same performance in terms of the $L_{2}$-error consistently with 1. Similarly, the same results hold for the standard deviation of the errors 5(b) 6(b), 7(b). Overall, as it is evident from all experiments 5, 6 and 7, after $M=100$ trees the centered KeRF and the simplified directional KeRF essentially coincide.

### v3

@@ 7. Proofs of theorems
We assume furthermore that all random variables are real-valued and $\lvert\lvert X\rvert\rvert_{L_{p}}\colon=(\mathbb{E}\lvert X\rvert^{p})^{\frac{1}{p}}$ and $\lvert\lvert X\rvert\rvert_{\infty}\colon=\inf\{t\colon P(\lvert X\rvert\leq t)=1\}$
We begin with this basic lemma providing tail bounds for centered iid random variables with bounded variance and supremum norm.
Lemma 1.
Let $X_{1},...,X_{n}$ be a sequence of real independent and identically distributed random variables with $\mathbb{E}(X_{i})=0.$ Assuming also that there is a uniform bound for the $L_{2}$-norm and the supremum norm i.e. $\mathbb{E}(\rvert X_{i}\lvert)^{2}\leq Ma_{n},$ $\lvert\lvert X_{i}\rvert\rvert_{\infty}\leq M\leq n$ for every $i=1,...,n.$ Then for every $t\leq 2\sqrt{Ma_{n}}$
($\mathbb{P}\big{(}\{\frac{\lvert\sum_{i=1}^{n}X_{i}\rvert}{n}\geq t\}\big{)}\leq 2\exp{(-\frac{t^{2}n}{Ma_{n}})}.$
Proof.
$\mathbb{P}\biggl{(}\frac{1}{n}\sum_{i=1}^{n}X_{i}\geq t\biggr{)}$ $=\mathbb{P}\biggl{(}\frac{\lambda}{n}\sum_{i=1}^{n}X_{i}\geq\lambda t\biggr{)}$
$=\mathbb{P}\biggl{(}\exp{(\frac{\lambda}{n}\sum_{i=1}^{n}X_{i})}\geq\exp{(\lambda t)\biggr{)}}$
$\leq\exp{(-\lambda t)}\mathbb{E}\exp\biggl{(}\frac{\lambda}{n}\sum_{i=1}^{n}X_{i}\biggl{)}$
$\leq\exp{(-\lambda t)}\prod_{i=1}^{n}\mathbb{E}\exp\biggl{(}\frac{\lambda}{n}X_{i}\biggl{)}.$
Where the inequalities are provided by Chebysev inequality and the independence of the random variables. Moreover, for convenience, let $Y_{j}=\frac{Xj}{n}$ and we observe that, $\lvert\lvert Y_{i}\rvert\rvert_{\infty}\leq 1$ and $\mathbb{E}(Y_{i})^{2}\leq\frac{Ma_{n}}{n^{2}}=\sigma^{2}$
$\mathbb{E}\exp\biggl{(}\frac{\lambda}{n}X_{i}\biggl{)}$ $=\mathbb{E}\biggl{(}1+\sum_{k=2}^{\infty}\frac{\lambda^{k}Y_{i}^{k}}{k!}\biggl{)}$
$=1+\sum_{k=2}^{\infty}\frac{\lambda^{k}\mathbb{E}(Y_{i}^{k})}{k!}$
$\leq 1+\sum_{k=2}^{\infty}\frac{\lambda^{k}(\mathbb{E}Y_{i}^{2})^{\frac{k}{2}}\lvert\lvert Y_{i}\rvert\rvert_{\infty}^{\frac{k}{2}}}{k!}$
$\leq 1+\sum_{k=2}^{\infty}\frac{\lambda^{k}((\sigma^{2})^{\frac{1}{2}})^{k}}{k!}$
$=1+\exp{(\lambda\sigma)}-\lambda\sigma-1$
$\leq 1+\lambda\sigma+(\lambda\sigma)^{2}-\lambda\sigma$
$\leq\exp{(\lambda\sigma)^{2}}$
where we have used that $\exp{(\lambda\sigma)}\leq 1+\lambda\sigma+(\lambda\sigma)^{2}$ when $\lambda\sigma\leq 1$ and $1+x\leq e^{x}.$ Hence,
$\mathbb{P}\biggl{(}\frac{1}{n}\sum_{i=1}^{n}X_{i}\geq t\biggr{)}$ $\leq\exp{(-\lambda t)}\exp{(\lambda\sigma)^{2}n}$
$\leq\exp{(-\frac{t^{2}}{2\sigma^{2}n})}$
$=\exp{(-\frac{t^{2}n}{Ma_{n}})}$
where we choose $\lambda=\frac{t}{2\sigma^{2}n}$ which is an accepted value of $\lambda\iff t\leq 2\sqrt{Ma_{n}}.$ We conclude the proof by replacing $X_{i}$ with $-X_{i}.$
Now we move to the next necessary lemma. Here our target random variables are multiplied by centered independent Gaussian noises, but we can still obtain a similar tail bound by getting advantage the shape of the Gaussian tales.
Lemma 2.
Let $X_{1},...,X_{n}$ be a non-negative sequence of independent and identically distributed random variables with $\mathbb{E}(X_{i})^{2}\leq Ma_{n},$ $\lvert\lvert X_{i}\rvert\rvert_{\infty}\leq M\leq n$ for every $i=1,...,n.$ Let also a sequence of independent random variables $\epsilon_{i}$ following normal distribution with zero mean and finite variance $\tilde{\sigma}^{2},$ for every $i=1,...,n.$ We assume also that $\epsilon_{i}$ are independent from $X_{i}$ for every $i=1,...,n.$
Then for every $t\leq 2\sqrt{Ma_{n}}$
($\mathbb{P}\big{(}\{\frac{\lvert\sum_{i=1}^{n}X_{i}\rvert}{n}\geq t\}\big{)}\leq 2\exp{(-\frac{t^{2}n}{Ma_{n}\tilde{\sigma}^{2}})}.$
Proof.
First of all, we observe from the proof of Lemma 1 that,
$\mathbb{P}\biggl{(}\frac{1}{n}\sum_{i=1}^{n}X_{i}\epsilon_{i}\geq t\biggr{)}\leq\exp{(-\lambda t)}\prod_{i=1}^{n}\mathbb{E}\exp\biggl{(}\frac{\lambda}{n}X_{i}\epsilon_{i}\biggl{)},$
and
$\mathbb{E}\exp\biggl{(}\frac{\lambda}{n}X_{i}\epsilon_{i}\biggl{)}$ $=\mathbb{E}\biggl{(}1+\sum_{k=2}^{\infty}\frac{\lambda^{k}Y_{i}^{k}\epsilon_{i}}{k!}\biggl{)}$
$\leq 1+\sum_{k=2}^{\infty}\frac{\lambda^{k}(\mathbb{E}Y_{i}^{2})^{\frac{k}{2}}\lvert\lvert Y_{i}\rvert\rvert_{\infty}^{\frac{k}{2}}\mathbb{E}(\epsilon_{i})^{k}}{k!}$
$\leq 1+\sum_{k=2}^{\infty}\frac{\lambda^{k}((\sigma^{2})^{\frac{1}{2}})^{k}\mathbb{E}(\epsilon_{i})^{k}}{k!}$
$=\mathbb{E}\exp\biggl{(}\lambda\sigma\epsilon_{i}\biggr{)}$
$\leq\mathbb{E}\exp\biggl{(}\lambda^{2}\sigma^{2}\tilde{\sigma}^{2}\biggr{)}$
Finally,
$\mathbb{P}\biggl{(}\frac{1}{n}\sum_{i=1}^{n}X_{i}\epsilon_{i}\geq t\biggr{)}$ $\leq\exp{(-\lambda t)}\exp{(\lambda\sigma\tilde{\sigma})^{2}n}$
$\leq\exp{(-\frac{t^{2}}{2\sigma^{2}\tilde{\sigma}^{2}n})}$
$=\exp{(-\frac{t^{2}n}{Ma_{n}})}$
where we choose $\lambda=\frac{t}{2\sigma^{2}\tilde{\sigma}^{2}n}$ which is an accepted value of $\lambda\iff t\leq 2\sqrt{Ma_{n}}.$ We conclude the proof by replacing $X_{i}$ with $-X_{i}.$
Theorem 6.
$\textbf{Y}=m(\textbf{X})+\epsilon$ where $\epsilon$ is a zero mean Gaussian noise with finite variance $\sigma$ independent of X. Assuming also that X is uniformly distributed in $[0,1]^{d}$ and $m$ is a Lipschitz function. Then there exists constants $c_{1},c_{2}$ depending on $d,\sigma$ and $\lVert m\rVert_{\infty}=\sup_{x\in[0,1]^{d}}|m(x)|$ such that,
$\mathbb{E}(\tilde{m}_{n,\infty}^{cc}(x)-m(x))^{2}\leq c_{1}\left(1-\frac{1}{2d}\right)^{2\log_{2}{n}}+C_{3}\frac{\log_{2}({\log_{2}n})^{d}}{(\log_{2}n)^{\frac{d-1}{2}}}\log{\bigg{(}\frac{\log{n}(\log_{2}n)^{\frac{d-1}{2}}}{\log_{2}({\log_{2}n})^{d}}\bigg{)}}$
Proof.
Following the notation in [30], let $x\in[0,1]^{d}$, $\lVert m\rVert_{\infty}=\sup_{x\in[0,1]^{d}}|m(x)|$, and by the construction of the algorithm
$\tilde{m}_{n,\infty}^{Cen}(x)=\frac{\sum_{i=1}^{n}Y_{i}K_{k}(x,X_{i})}{\sum_{i=1}^{n}K_{k}(x,X_{i})}.$
Let
$A_{n}(x)=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{Y_{i}K_{k}(x,X_{i})-\mathbb{E}(YK_{k}(x,X))}{\mathbb{E}(K_{k}(x,X))}\right),$
$B_{n}(x)=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{K_{k}(x,X_{i})-\mathbb{E}(K_{k}(x,X))}{\mathbb{E}(K_{k}(x,X))}\right),$
and
$M_{n}(x)=\frac{\mathbb{E}(YK_{k}(x,X))}{\mathbb{E}(K_{k}(x,X))}.$
Hence, we can reformulate the estimator as
$\tilde{m}_{n,\infty}^{Cen}(x)=\frac{M_{n}(x)+A_{n}(x)}{B_{n}(x)+1}.$
Let $t\in(0,\frac{1}{2})$ and the event $C_{t}(x)$ where $\{A_{n}(x),B_{n}(x)\leq t\}.$
$\mathbb{E}(\tilde{m}_{n,\infty}^{cc}(x)-m(x))^{2}$ $=\mathbb{E}(\tilde{m}_{n,\infty}^{cc}(x)-m(x))^{2}\mathbbm{1}_{C_{t}(x)}+\mathbb{E}(\tilde{m}_{n,\infty}^{cc}(x)-m(x))^{2}\mathbbm{1}_{C_{t}^{c}(x)}$
$\leq\mathbb{E}(\tilde{m}_{n,\infty}^{cc}(x)-m(x))^{2}\mathbbm{1}_{C_{t}^{c}(x)}+c_{1}\left(1-\frac{1}{2d}\right)^{2k}+c_{2}t^{2}.\quad$
Where the last inequality was obtained in [30, p.1496] Moreover, in [30],
$\mathbb{E}(\tilde{m}_{n,\infty}^{cc}(x)-m(x))^{2}\mathbbm{1}_{C_{t}^{c}(x)}\leq c_{3}(\log{n})(\mathbb{P}(C_{t}^{c}(x)))^{\frac{1}{2}}.$
Proposition 4.
The probability $\mathbb{P}(C_{t}^{c}(x))\leq C\exp{(-\frac{t^{2}n}{2^{k}a_{n}})}$ for a constant $C$ independent of $k,n$ and $a_{n}$ is a sequence that converges to zero as $n$ tends to infinity.
Proof.
First of all, we notice that
$\mathbb{P}(C_{t}^{c}(x))\leq\mathbb{P}(\rvert A_{n}(x)\rvert>t)+\mathbb{P}(\lvert B_{n}(x)\rvert>t).$
The related result will follow by working separately on both inequalities.
By lemma B.4 [2] for all $x\in[0,1]^{d}$ and $d\geq 2$ we have
$\mathbb{E}\big{(}K^{c}_{k}(x,X)\big{)}^{2}\leq\frac{C_{1}+C_{2}(\log_{2}(k))^{d}}{k^{\frac{d-1}{2}}2^{k}}$
where $C_{1}=1+\frac{2d^{\frac{d}{2}}}{(4\pi)^{\frac{d-1}{2}}}$ , $C_{2}=5^{d}(\frac{d-1}{2})^{d}$ and for convenience let $a_{n}=\frac{C_{1}+C_{2}(\log_{2}(k))^{d}}{k^{\frac{d-1}{2}}}.$
Hence, let $\tilde{X}_{i}=\frac{K_{k}(x,X_{i})}{\mathbb{E}(K_{k}(x,X))}-1$ a sequence of centered iid random variables with
$\lvert\lvert\tilde{X}_{i}\rvert\rvert_{\infty}=\sup\{\lvert\tilde{X}_{i}\rvert\}=\sup\{\lvert\frac{K_{k}(x,X_{i})}{\mathbb{E}(K_{k}(x,X))}-1\rvert\}\leq\frac{1}{\mathbb{E}(K_{k}(x,X))}\sup{K_{k}(x,X_{i})}+1\leq 2^{k+1}$
and
$\mathbb{E}\big{(}\frac{K^{c}_{k}(x,X_{i})}{\mathbb{E}(K^{c}_{k}(x,X)}\big{)}^{2}$ $=\frac{1}{(\mathbb{E}(K^{c}_{k}(x,X)))^{2}}\mathbb{E}\big{(}K^{c}_{k}(x,X)\big{)}^{2}$
$\leq\frac{1}{(\frac{1}{2^{k}})^{2}}\frac{C_{1}+C_{2}(\log_{2}(k))^{d}}{k^{\frac{d-1}{2}}2^{k}}$
$=2^{k}a_{n}$
By lemma 1, for every $t\leq 2\sqrt{2^{k}a_{n}}$
$\mathbb{P}(\lvert B_{n}(x)\rvert>t)\leq 2\exp{(-C\frac{t^{2}n}{2^{k}a_{n}})}$
We need an estimate for the $\mathbb{P}(\rvert A_{n}(x)\rvert>t)$ where,
$A_{n}(x)=\frac{1}{n}\sum_{i=1}^{n}\big{(}\frac{Y_{i}K_{k}(x,X_{i})-\mathbb{E}(YK_{k}(x,X))}{\mathbb{E}(K_{k}(x,X))}\big{)}.$
With simple calculations,
$A_{n}(x)$ $=\frac{1}{n}\sum_{i=1}^{n}\bigg{(}\frac{Y_{i}K_{k}(x,X_{i})-\mathbb{E}(YK_{k}(x,X))}{\mathbb{E}(K_{k}(x,X))}\bigg{)}$
$=\frac{1}{n}\sum_{i=1}^{n}\bigg{(}\frac{m(X_{i})K_{k}(x,X_{i})-\mathbb{E}(m(X)K_{k}(x,X))}{\mathbb{E}(K_{k}(x,X))}\bigg{)}+\frac{1}{n}\sum_{i=1}^{n}\bigg{(}\frac{\epsilon_{i}K_{k}(x,X_{i})-\mathbb{E}(\epsilon K_{k}(x,X))}{\mathbb{E}(K_{k}(x,X))}\bigg{)}$
$=\frac{1}{n}\sum_{i=1}^{n}\bigg{(}\frac{m(X_{i})K_{k}(x,X_{i})-\mathbb{E}(m(X)K_{k}(x,X))}{\mathbb{E}(K_{k}(x,X))}\bigg{)}+\frac{1}{n}\sum_{i=1}^{n}\bigg{(}\frac{\epsilon_{i}K_{k}(x,X_{i})}{\mathbb{E}(K_{k}(x,X))}\bigg{)}.$
Let $Z_{i}=2\frac{\epsilon_{i}K_{k}(x,X_{i})}{\mathbb{E}(K_{k}(x,X))}$ a sequence of centered iid random variables with
$\mathbb{E}(Z_{i})^{2}$ $\leq\frac{4}{\frac{1}{(2^{k})^{2}}}\mathbb{E}(\epsilon K_{k}(x,X))^{2}$
$=\frac{4}{\frac{1}{(2^{k})^{2}}}\mathbb{E}(\epsilon)^{2}\mathbb{E}(K_{k}(x,X)))^{2})$
$\leq\tilde{c}2^{k}\sigma^{2}a_{n}$
and
$\lvert\lvert\frac{K_{k}(x,X_{i})}{\mathbb{E}(K_{k}(x,X))}\rvert\rvert_{\infty}\leq 2^{k}$
hence by 2
$\mathbb{P}\big{(}\lvert Z_{i}\rvert\geq t\big{)}=\mathbb{P}\big{(}\frac{2}{n}\sum_{i=1}^{n}\lvert\frac{\epsilon_{i}K_{k}(x,X_{i})}{\mathbb{E}(K_{k}(x,X))}\rvert\geq t\big{)}\leq 2\exp{(-\frac{t^{2}n}{2^{k}a_{n}}})$
for every $t\leq 2\sqrt{2^{k}a_{n}}$
$\mathbb{P}\big{(}\lvert A_{n}(x)\rvert\geq t\big{)}$ $\leq\mathbb{P}\bigg{(}\bigg{\lvert}\frac{2}{n}\sum_{i=1}^{n}\frac{m(X_{i})K_{k}(x,X_{i})-\mathbb{E}(m(X)K_{k}(x,X))}{\mathbb{E}(K_{k}(x,X))}\bigg{\rvert}\geq t\bigg{)}$
$\quad+\mathbb{P}\bigg{(}\bigg{\lvert}\frac{2}{n}\sum_{i=1}^{n}\frac{\epsilon_{i}K_{k}(x,X_{i})}{\mathbb{E}(K_{k}(x,X))}\bigg{\rvert}\geq t\bigg{)}$
$\leq 2\exp{(-C_{1}\frac{t^{2}n}{2^{k}}})+2\exp{(-C_{2}\frac{t^{2}n}{2^{k}a_{n}}})\leq C\exp{(-C_{3}\frac{t^{2}n}{2^{k}a_{n}}}).$
where we have used a partial result from [21] [proposition 6.] and finally,
$\mathbb{P}(C_{t}^{c}(x))\leq 2\exp{(-\tilde{C}\frac{t^{2}n}{2^{k}a_{n}})}.$
To obtain the desired rate of convergence, we need an upper bound for
$\mathbb{E}(\tilde{m}_{n,\infty}^{cc}(x)-m(x))^{2}\leq c_{3}\log{n}\exp{(-\tilde{C}\frac{t^{2}n}{2^{k}a_{n}})}+c_{1}\left(1-\frac{1}{2d}\right)^{2k}+c_{2}t^{2}$
and we choose $2^{k}=n$ in the mean interpolation regime,
$\mathbb{E}(\tilde{m}_{n,\infty}^{cc}(x)-m(x))^{2}\leq c_{3}\log{n}\exp{(-\tilde{C}\frac{t^{2}}{a_{n}})}+c_{1}\left(1-\frac{1}{2d}\right)^{2\log_{2}{n}}+c_{2}t^{2}$
Finally, by minimizing the right hand of the equation in terms of $t$ one has that, $t^{2}=Ca_{n}\log{\frac{\log{n}}{a_{n}}}$ and of course,
$c_{3}\log{n}\exp{(-\tilde{C}\frac{t^{2}}{a_{n}})}+c_{1}\left(1-\frac{1}{2d}\right)^{2\log_{2}{n}}+c_{2}t^{2}$ $\leq c_{1}\left(1-\frac{1}{2d}\right)^{2\log_{2}{n}}$
$\quad+C_{3}a_{n}+C_{2}a_{n}\log{(\frac{\log{n}}{a_{n}})}$
and therefore, since $a_{n}=\frac{\log_{2}({\log_{2}n})^{d}}{(\log_{2}n)^{\frac{d-1}{2}}}$
$\mathbb{E}(\tilde{m}_{n,\infty}^{cc}(x)-m(x))^{2}\leq c_{1}\left(1-\frac{1}{2d}\right)^{2\log_{2}{n}}+C_{3}\frac{\log_{2}({\log_{2}n})^{d}}{(\log_{2}n)^{\frac{d-1}{2}}}\log{\bigg{(}\frac{\log{n}(\log_{2}n)^{\frac{d-1}{2}}}{\log_{2}({\log_{2}n})^{d}}\bigg{)}}$
and this concludes the proof. ∎